let A and B be two nonsingular matrices such that A is not equal to I,B3=I and AB = BA2, find the least value k such that A raise to the power k is equal to I.

Grade:12

3 Answers

Amit Jain
10 years ago
Please re-post the question as it is not clear what is asked(the powers and coefficients.)
akash
10 years ago
let A and B be two nonsingular matrices such that A is not equal to I,B^3=I and AB = BA^2, find the least value k such that A^k=I.
Testing
10 years ago
Dear Student, The answer for this post is sent in a separate attachment through email. Thanks, Askiitians Support Team